a car is traveling on a road that passed through rolling terrain. at the bottom of one particular dip in the road which may be approximately as a vertical circle of radius 65m, the car and driver are moving with an instantaneous horizontal velocity of 17m/s. if the mass of the driver is 80kg what is the apparent weight of the driver at the bottom of the dip
1
Expert's answer
2019-11-04T10:40:45-0500

Apparent weight = "mg-\frac{mv^2}{r}"

= "(80\times10)-(\frac{80\times17^2}{65})=800-355.70=444.3\ N"